Operations Report – December 16th

Crew 201 Operations Report 16-Dec-2018

SOL:1

Name of person filing report: César Augusto Serrano Baza

Non-nominal systems: Generator Autostart

Notes on non-nominal systems: None
Generator (hours run): Around 5:30 last night to around 9 am today.

Solar— SOC 81%
Diesel Reading – 90%
Propane Reading – 92%
Ethanol Free Gasoline – N/A gallons.

Water (Axillary tank) – 0 gallons
Water (static tank) – 340 gallons (Refilled today)
Axillary to Static tank transfer– no
Gallons transferred: 0
Water in GreenHab – High level (refilled today)
Water (loft) – Static to Loft Pump used – yes
Water Meter: 5.2

Toilet tank emptied: NO

Deimos rover used: no, still not functional
Hours: 0
Beginning charge: –
Ending charge: –
Currently charging: no

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: 0.83 hrs
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Commander and XO
Hours:0.83
Beginning charge: 17:31
Ending charge: 100%
Currently charging: YES

Opportunity rover used: Engineer and GreenHab Officer
Hours: 0.83
Beginning charge: 17:31
Ending charge: 100%
Currently charging: YES

Curiosity rover used: Astronomer and HSO
Hours: 0.83
Beginning charge: 17:32
Ending charge: 100%
Currently charging: YES

Notes on rovers: Rovers where used today for the crew trainning.

# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now parked away from the station.

HabCar used and why, where? Yes, used for refill the static tank water and the GreenHab water (1 time each tank) to Hanksville gas station.
CrewCar used and why, where? Last crew took it wth them.

General notes and comments: We said goodbye to crew 200. We installed a heather in the Dome.

Summary of internet: Nothing to report.

Summary of suits and radios: We had a quick trainning about how to wear the space suit.

Summary of Hab operations: Nothing to report.

Summary of GreenHab operations: Keep water in the plants.

Summary of ScienceDome operations: Nothing to report.

Summary of RAMM operations: Nothing to report.

Summary of any observatory issues: Musk Solar observatory – Nominal. RObotic Observatory – In use at 17:30 hrs.

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A